Given the demand for a product as Qd = 50 - 8P and the supply is given as Qs = -17,5 + 10P. You are told equilibrium is obtained at the point where Qd = Qs. The equilibrium quantity for the product is

1
Expert's answer
2021-06-11T11:47:07-0400

Let's first find the equilibrium price for the product:


"Q_d=Q_s,""50 - 8P_E=-17,5 + 10P_E,""67.5=18P_E,""P_E=\$3.75."

Finally, we can find the equilibrium quantity for the product by substituting the equilibrium price either in the demand or supply function:


"Q_E=50-\$3.75\cdot8=20."